OpenStudy (anonymous):

What is the 9th term of the geometric sequence where a1 = 1024 and a3 = 64?

OpenStudy (anonymous):

\[a_1,a _2=a_1r, a_3=a_1r^2, ...\] so you have \[\frac{a_3}{a_1}=r^2\] and therefore \[\frac{1024}{64}=r^2\] \[16=r^2\]solve for \(r\)

OpenStudy (mathmagician):

1024*k^2=64 k^2=64/1024=1/16 k=1/4 therefore 9th term is 1024/(4^8)=1024/65536=1/64

OpenStudy (anonymous):

lol i had it upside down! \[\frac{64}{1024}=r^2\] \[\frac{1}{16}=r^2\] \[r=\frac{1}{4}\]

OpenStudy (anonymous):

it either 2,028 1,728 1,452 2,268

OpenStudy (anonymous):

actually it is none of these

OpenStudy (anonymous):

\[a, ar, ar^2, ar^3, ar^4, ...\] \[1024,256,64,16,...\] it is a decreasing sequence